OpenStudy (anonymous):

A gaseous hydrogen and carbon containing compound is decomposed and found to contain 82.66% carbon and 17.34% hydrogen by mass. The mass of 158 mL of the gas, measured at 556 mmHg and 25 ∘C, was found to be 0.275 g. What is the molecular formula of the compound?

OpenStudy (aaronq):

use the ideal gas law and rearrange it for molecular mass, M. PV=nRT -> PV=mRT/M -> M=mRT/PV plug in your values, find M find the mass that contributes to each of the elements (multiply M by their %) and find how many of atoms of each are present in the formula.

OpenStudy (anonymous):

I'm still lost.

OpenStudy (anonymous):

ok i tried the ideal gas law for M of carbon, and i got a really high number (1440). not sure if i'm doing it right

OpenStudy (anonymous):

I think you tried it as i did because i got the same number but it said it was incorrect. Thank you :)

OpenStudy (anonymous):

the answer is Carbon- 1 Hydrogen-2 Fluorine- 2

OpenStudy (anonymous):

@kishnktty: Let UNITs guide you; always USE THEM in your calculation to prevent errors moles = PV/RT = (556 mm/760mm/atm)*0.158liter / (0.08205liter-atm/moleK * 298K) = ?? MW = 0.275 g / ?? moles = ?? Empirical formula = C(82.66/12.01)H(17.34/1.008) ---- C1H2.5 MW/Empirical W ==> Molecular Formula/Empirical Formula

OpenStudy (anonymous):

ok i got most of what you're saying, @save. but how can i find the molecular formula using the MW,Emp W, and Emp formula?

OpenStudy (aaronq):

\(M=\dfrac{(0.275 g)*(0.08206)*(25+273)}{(\dfrac{556}{760} atm)*(0.158 L)}=58.1784 ~g/mol\) C = 82.66%*58.1784 g/mol=48.1 g/mol/12 g/mol= 4 Carbon atoms H=17.34%*58.1784 g/mol= 10.1 g/mol/1 /mol = 10 Hydrogen atoms \(C_4H_{10}\) which is butane

OpenStudy (aaronq):

y'all did it wrong